# Charm quark evolution in the early stages of heavy-ion collisions

### v1

@@ I Introduction
Heavy-ion collision experiments conducted at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) offer a unique opportunity to study QCD matter under extreme but controlled conditions. Model-to-data comparisons are the main tool for inferring the properties of the QCD matter produced. However, modeling heavy-ion collisions is complex and requires a sophisticated multistage framework. The current state-of-the-art approach incorporates initial states with information about the nucleonic and sub-nucleonic degrees of freedom [1, 2], hydrodynamical evolution of the Quark-Gluon Plasma (QGP) [3, 4, 5], and hadronic transport [6, 7]. The multi-stage model parameters can be constrained using experimental data by performing Bayesian analyses [8, 9, 10, 11].
@@
In this study, we utilize the IP-Glasma model [2, 39] coupled to the viscous hydrodynamic simulation MUSIC [40, 41, 42] and hadronic transport UrQMD [6], to model the bulk medium evolution. This framework has been successful in describing a wide range of soft hadronic observables in heavy-ion collisions [43]. The model is based on describing the nuclei, prior to their collision, as two sheets of color glass condensate, which captures the essence of the low momentum-fraction gluons present in the nuclei accelerated to high energies. The gluon field generated by the collision of the energetic nuclei evolves for a short time according to the classical Yang-Mills equations. The energy-momentum tensor of these gluons then serves as the input for viscous hydrodynamics, which takes over the time evolution of the bulk medium. We simulate the evolution of charm quarks starting from the fluctuating IP-Glasma initial stage, followed by their evolution in the hydrodynamically expanding QGP, using the MARTINI event generator [44, 45]. By incorporating the charm quark dynamics in the initial stage and in the QGP phase, we provide first phenomenological insights into charm quark evolution within the IP-Glasma initial state and its effects on the final state observables.
@@ II.1 Bulk medium evolution
with $J^{a\,\mu}$ the current term that sources soft gluons and $F^{a}_{\mu\nu}$ the field strength tensor, where $a$ denotes the color index. The evolution of the gluon fields is governed by the Classical Yang-Mills (CYM) equations:
@@
where $\theta=\partial_{\mu}u^{\mu}$ is the scalar expansion rate, ${\pi}^{\langle\mu\nu\rangle}=\Delta^{\mu\nu}_{\alpha\beta}\pi^{\alpha\beta}$, and $\sigma^{\mu\nu}=\Delta^{\mu\nu}_{\alpha\beta}\partial^{\alpha}u^{\beta}$ with the symmetric, traceless projection operator defined as $\Delta^{\mu\nu}_{\alpha\beta}\equiv\frac{1}{2}(\Delta^{\mu}_{\alpha}\Delta^{\nu}_{\beta}+\Delta^{\mu}_{\beta}\Delta^{\nu}_{\alpha})-\frac{1}{3}\Delta^{\mu\nu}\Delta_{\alpha\beta}$. This framework is implemented numerically in MUSIC. The first-order viscous coefficients, the shear and bulk viscosities, are tuned to data in [43]. The second-order transport coefficients $\delta_{\pi\pi},\phi_{7},\tau_{\pi\pi},\lambda_{\pi\Pi},\tau_{\pi},\delta_{\Pi\Pi},\lambda_{\Pi\pi},\tau_{\Pi}$ are expressed in terms of shear and bulk viscosities [49].
@@ II.2 Heavy quark production
We implement a formation time of $\tau_{F}=1/m_{c}$ in the charm quark’s rest frame. Charm quarks are allowed to interact with the medium only after their formation time has passed. Before the formation time is reached, charm quarks/anti-quarks are bound to their original partner via a Cornell potential [55]. We will study the effect of varying the formation time below. Charm quark positions in the background medium at midrapidity for a single event are shown in Fig. 1.
@@ II.4 Heavy quark hadronization
where the subscript $M$ denotes mesons, $N_{coal}$ is a normalization factor, and $g_{M}$ is the degeneracy factor of the mesons. The width parameter of the Gaussian distribution is defined as $\sigma=1/\sqrt{\mu\omega}$ where $\mu=m_{c}m_{u/d}/(m_{c}+m_{u,d})$ is the reduced mass of the heavy and light quark system. For quantitative estimation, we consider the thermal mass of light quarks as $m_{u,d}=300$ MeV. The angular frequency of the oscillator $\omega$ is taken to be $0.106$ GeV, as tuned in Refs. [18, 66]. The relative momentum in the two-body center-of-mass frame is
@@ Acknowledgments
This research used the resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC awards NP-ERCAP0033578 and NP-ERCAP0032155. This work is supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under DOE Contract No. DE-SC0024711 (M.S.) and DE-SC0012704 (B.P.S.) and within the framework of the Saturated Glue (SURGE) Topical Theory Collaboration (B.P.S.). M.S. is also supported by the Vanderbilt University. M.K. acknowledges the Special Postdoctoral Researchers Program of RIKEN, the Faculty Research Scheme (FRS project number: MISC 0240) at IIT (ISM) Dhanbad, and the Department of Science and Technology (DST), Govt. of India, for the INSPIRE-Faculty award (DST/INSPIRE/04/2024/001794). S.J. and C.G. are supported by the Natural Sciences and Engineering Research Council of Canada under grant numbers SAPIN-2018-00024 and SAPIN-2020-00048, respectively.

### v2

@@ I Introduction
Heavy-ion collision experiments conducted at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) offer a unique opportunity to study QCD matter under extreme but controlled conditions. Model-to-data comparisons are the main tool for inferring the properties of the Quantum Chromodynamics (QCD) matter produced. However, modeling heavy-ion collisions is complex and requires a sophisticated multistage framework. The current state-of-the-art approach incorporates initial states with information about the nucleonic and sub-nucleonic degrees of freedom [1, 2], hydrodynamical evolution of the Quark-Gluon Plasma (QGP) [3, 4, 5], and hadronic transport [6, 7]. The multi-stage model parameters can be constrained using experimental data by performing Bayesian analyses [8, 9, 10, 11].
@@
In this study, we utilize the impact parameter dependent glasma (IP-Glasma) model [2, 39] coupled to the viscous hydrodynamic simulation MUSIC (which is short for ‘MUSCL for Ion Collisions’, where MUSCL stands for Monotonic Upstream-centered Scheme for Conservation Laws) [40, 41, 42] and the hadronic transport model Ultrarelativistic Quantum Molecular Dynamics (UrQMD) [6], to model the bulk medium evolution. This framework has been successful in describing a wide range of soft hadronic observables in heavy-ion collisions [43]. The model is based on describing the nuclei, prior to their collision, as two sheets of color glass condensate, which captures the essence of the low momentum-fraction gluons present in the nuclei accelerated to high energies. The gluon field generated by the collision of the energetic nuclei evolves for a short time according to the classical Yang-Mills equations. The energy-momentum tensor of these gluons then serves as the input for viscous hydrodynamics, which takes over the time evolution of the bulk medium. We simulate the evolution of charm quarks starting from the fluctuating IP-Glasma initial stage, followed by their evolution in the hydrodynamically expanding QGP, using the MARTINI (Modular Algorithm for Relativistic Treatment of Heavy Ion Interactions) event generator [44, 45]. By incorporating the charm quark dynamics in the initial stage and in the QGP phase, we provide first phenomenological insights into charm quark evolution within the IP-Glasma initial state and its effects on the final state observables.
@@ II.1 Bulk medium evolution
with $J^{a\,\mu}$ the current term that sources soft gluons field $A^{a}_{\mu}$, and $F^{a}_{\mu\nu}$ the field strength tensor, where $a$ denotes the color index. The evolution of the gluon fields is governed by the Classical Yang-Mills (CYM) equations:
@@
where $\theta=\partial_{\mu}u^{\mu}$ is the scalar expansion rate, ${\pi}^{\langle\mu\nu\rangle}=\Delta^{\mu\nu}_{\alpha\beta}\pi^{\alpha\beta}$, and $\sigma^{\mu\nu}=\Delta^{\mu\nu}_{\alpha\beta}\partial^{\alpha}u^{\beta}$ with the symmetric, traceless projection operator defined as $\Delta^{\mu\nu}_{\alpha\beta}\equiv\frac{1}{2}(\Delta^{\mu}_{\alpha}\Delta^{\nu}_{\beta}+\Delta^{\mu}_{\beta}\Delta^{\nu}_{\alpha})-\frac{1}{3}\Delta^{\mu\nu}\Delta_{\alpha\beta}$. This framework is implemented numerically in MUSIC. The first-order viscous coefficients, the shear ($\eta$) and bulk ($\zeta$) viscosities, are tuned to data in [43]. The second-order transport coefficients $\delta_{\pi\pi},\phi_{7},\tau_{\pi\pi},\lambda_{\pi\Pi},\tau_{\pi},\delta_{\Pi\Pi},\lambda_{\Pi\pi},\tau_{\Pi}$ are expressed in terms of shear and bulk viscosities [49].
@@ II.2 Heavy quark production
We implement a formation time of $\tau_{F}=1/m_{c}$ in the charm quark’s rest frame. Here $m_{c}$ is the mass of a charm quark. Charm quarks are allowed to interact with the medium only after their formation time has passed. Before the formation time is reached, charm quarks/anti-quarks are bound to their original partner via a Cornell potential [55]. We will study the effect of varying the formation time below. Charm quark positions in the background medium at midrapidity for a single event are shown in Fig. 1.
@@ II.4 Heavy quark hadronization
where the subscript $M$ denotes mesons, $N_{coal}$ is a normalization factor, and $g_{M}$ is the degeneracy factor of the mesons. The quark distributions is assumed to be uniform in the volume $V$. The width parameter of the Gaussian distribution is defined as $\sigma=1/\sqrt{\mu\omega}$ where $\mu=m_{c}m_{u/d}/(m_{c}+m_{u/d})$ is the reduced mass of the heavy and light quark system. For quantitative estimation, we consider the thermal mass of light quarks as $m_{u/d}=300$ MeV. The angular frequency of the oscillator $\omega$ is taken to be $0.106$ GeV, as tuned in Refs. [18, 66]. The relative momentum in the two-body center-of-mass frame is
@@ IV Summary and Outlook
Data Availability
The data that support the findings of this article are openly available [76].
@@ Acknowledgments
This research used the resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC awards NP-ERCAP0033578 and NP-ERCAP0032155. This work is supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under DOE Contract No. DE-SC0024347 (M.S.) and DE-SC0012704 (B.P.S.) and within the framework of the Saturated Glue (SURGE) Topical Theory Collaboration (B.P.S.). M.S. is also supported by the Vanderbilt University. M.K. acknowledges the Special Postdoctoral Researchers Program of RIKEN, the Faculty Research Scheme (FRS project number: MISC 0240) at IIT (ISM) Dhanbad, and the Department of Science and Technology (DST), Govt. of India, for the INSPIRE-Faculty award (DST/INSPIRE/04/2024/001794). S.J. and C.G. are supported by the Natural Sciences and Engineering Research Council of Canada under grant numbers SAPIN-2024-00026 and SAPIN-2020-00048, respectively.